A computer program selects an integer in the set \(\{k : 1 \le k \le 10,00,000\}\) at random and prints out the result. This process is repeated 1 million times. What is the probability that the value k = 1 appears in the printout atleast once ?
0.632121
Let \(n = 10^6\). In one selection, the probability that k = 1 is picked is \(\frac{1}{n}\), so the probability it is not picked is \(1 - \frac{1}{n}\).
The n selections are independent, so the probability that k = 1 never appears is \(\left(1 - \frac{1}{n}\right)^{n}\).
For large n, \(\left(1 - \frac{1}{n}\right)^{n} \approx e^{-1} \approx 0.367879\).
By the complement rule, P(at least once) \(= 1 - e^{-1} \approx 1 - 0.367879 = 0.632121\).
Hence, the answer is 0.632121.
Consider an experiment of tossing two fair dice, one black and one red. What is the probability that the number on the black die divides the number on red die ?
Match List - I with List - II.
| List - I (Queries) | List - II (Probability) |
|---|---|
| (A) A bag contains 6 white and 4 red balls. Two balls are drawn at random. What is the chance, they will be the same colour ? | (I) 3/68 |
| (B) In a pack of 52 cards, one card is drawn at random, what is the probability that it is either a king or a queen ? | (II) 14/68 |
| (C) A bag contains 6 red, 4 white and 8 blue balls. If three balls are drawn at random, find the probability of 1 red and 2 white balls ? | (III) 2/13 |
| (D) A bag contains 6 red, 4 white and 8 blue balls. If three balls are drawn at random. Find the probability of 2 blue and 1 red balls ? | (IV) 7/15 |
Choose the correct answer from the options given below :
A coin is tossed successively three times. Find the Probability (P), Event (E), Sample space (S) of getting exactly one head or two heads, where n is number of occurrence.
(A) n(S) = 8 and n(E) = 4
(B) n(E) = 6 and n(S) = 8
(C) P(E) = \(\dfrac{3}{4}\)
(D) P(E) = \(\dfrac{1}{2}\)
Choose the correct answer from the options given below :
The probability that A hits a target is \(\frac{1}{4}\), and the probability that B hits the target is \(\frac{2}{5}\). Both shoot at the target, what is the probability that at least one of them hits the target, i.e., that A or B (or both) hit the target?
Events A, Band C are mutually exclusive events such that \(P(A) = \dfrac{3x + 1}{3}, P(B) = \dfrac{1-x}{4}\)and \(P(C) = \dfrac{1-2x}{4}\)The set of possible values of x are in the interval
Let A, B be two events in a discrete probability space with ℙ(A) > 0 and ℙ(B) > 0. Which of the following are necessarily true?
For a distribution, the mean is 10, variance is 16, γ 1is + 1 and β 2is 4. The distribution is:
For the discrete distribution, the Pearson's coefficient of skewness β 2is always:
A six-face fair dice is rolled a large number of times. The mean value of the outcomes is ________